OpenStudy (anonymous):

I GIVE GOLD MEDAL! what is the value of f[g(3)] for the functions f(x) = x+2 and g(x) = 5x-1?

OpenStudy (anonymous):

f(g(3)) means to FIRST compute g(3), then whatever the answer is, compute f(of that answer). So, g(3) = 5(3) - 1 = 14 So, f(14) = 14 + 2 = 16

OpenStudy (anonymous):

Final answer is 16.

OpenStudy (anonymous):

thank you! @Easyaspi314